Circular No. 8125

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 J 19 AND S/2003 J 20 Further to IAUC 8116, the discoveries of two new satellites of Jupiter

have been reported by teams led by B. Gladman (S/2003 J 19) and by S. S. Sheppard (S/2003 J 20). Astrometry, preliminary orbital elements by B. G. Marsden, and ephemerides appear on MPEC 2003-G64 (S/2003 J 19, e = 0.33, i = 163^{o}, P = 701 days, H = 16.7) and MPEC 2003-G67 (S/2003 J 20, e = 0.29, i = 55^{o}, P = 455 days, H = 15.6).

SUPERNOVAE 2003dx, 2003dy, 2003dz, 2003ea, 2003eb A. Riess, Space Telescope Science Institute, reports [on behalf of the

GOODS Treasury Team (cf. IAUC 8052 and 8081, plus D. de Mello) and the Hubble Higher-z Supernova Team] the discovery of five apparent supernovae in images of the region around the Hubble Space Telescope ‘Deep Field’, each detected by subtracting ACS images taken on Feb. 20–23 from those taken on Apr. 4–5. None of these transients was detected in F850lp images taken on 2002 Nov. 21–25 and 2003 Jan. 2–6 (limiting mag 27.0). Each apparent supernova was detected independently in each of four separate z-band (F850lp) and two i-band (F775W) images. The morphology of each transient is consistent with that of the ACS Point Spread Function (FWHM 0′′.1). The lack of detectable motion (< 0′′.05) relative to other sources in the images over ∼ 45 min eliminates the possibility that these are small bodies in the solar system; other possibilites such as flare stars, closer cataclysmic variable stars, and active galactic nuclei are also dismissed as unlikely. The offsets given below are from the centers of the nearest galaxies visible in the ACS images.

SN  2003 UT  α_{2000}  δ_{2000}  z  Offset
2003dx Apr. 4.52 12^{h}36^{m}31.^{s}75 +62^{o}08′48′′.7 25.2 0′′.15 E, 0′′.10 N
2003dy Apr. 4.67 12 37 09.16 +62 11 29.0 24.5 0′′.25 E, 0′′.35 S
2003dz Apr. 4.79 12 36 39.91 +62 07 52.7 25.1 0′′.25 W
2003ea Apr. 5.65 12 37 12.04 +62 12 38.3 25.4 0′′.15 N
2003eb Apr. 5.65 12 37 15.18 +62 13 34.6 22.8 0′′.50 E, 0′′.75 S

COMET C/2003 H2 (LINEAR) Further to IAUC 8122, J. McGaha (Tucson, AZ) reports that six

stacked 2-min CCD exposures taken on Apr. 25.3 UT (0.30-m reflector) show a 6′′ coma and a 10′′ tail in p.a. 50^{o}.

2003 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green